# Polarimeter to Unify the Corona and Heliosphere (PUNCH)

### v1

@@ 1 Introduction
Joint studies of the solar corona and solar wind have generally required spanning the two radically different experimental disciplines, but have borne fruit in connecting the Sun and near-Earth environments (Zirker, 1977; Sheeley et al., 1985; Domingo, Fleck, and Poland, 1995). ESA’s Solar Orbiter mission (Müller et al., 2020) carries a hybrid instrument suite specifically designed to enable joint remote/in-situ studies of solar wind features and streams, by associating particular remotely sensed coronal features with the plasma detected in-situ by the spacecraft. Recently, the Parker Solar Probe (Fox et al., 2016) has carried in-situ sampling devices into the outer reaches of the solar corona itself (Kasper et al., 2021), allowing direct measurement of the microphysics in the outer corona with simultaneous remote sensing of the relevant large-scale structures surrounding the system under study.
@@
Multiple instruments, including the Coriolis Solar Mass Ejection Imager (Coriolis/SMEI; Eyles et al., 2003), the Solar-Terrestrial Relations Observatory’s Sun-Earth Connection Coronal and Heliospheric Investigation (STEREO/SECCHI; Howard et al., 2008), the Solar Orbiter Heliospheric Imager (SOLO/SOLOHI; Howard et al., 2020), and the Parker Solar Probe Wide-field Imager for Solar Probe (PSP/WISPR; Vourlidas et al., 2016), have been built to image large and bright features in the solar wind. With the advent of imaging from off the Sun-Earth line, via STEREO/SECCHI, direct, quantitative, simultaneous imaging and in-situ studies of Earth-bound events became possible (e.g., Davies et al., 2009; DeForest, Howard, and McComas, 2013), and the presence of multiple deep-space missions with intersecting fields of view and on-board in-situ instruments has greatly increased the number of such studies. PUNCH brings routine global imaging of the solar wind and its structures into the mix of available observations, explicitly enabling a unified treatment of the corona and solar wind as a single system.
@@
PUNCH polarimetry is designed to allow location of features in three dimensions (3D) from a single vantage point: both mesoscale solar wind structure (e.g., Viall, DeForest, and Kepko, 2021) and coronal mass ejections (de Koning, Pizzo, and Biesecker, 2009; Mierla et al., 2009; DeForest, de Koning, and Elliott, 2017). This is possible because the K corona consists of sunlight that is Thomson scattered by free electrons. Thomson scattering polarizes the outbound beam according to the scattering angle (Billings, 1966; Tappin and Howard, 2009; DeForest, Howard, and Tappin, 2013). Each coronal feature is therefore polarized to a degree that depends on its three-dimensional location. The basic theory has been known for decades, but was only reduced to practice once image quality and analysis methods improved enough to reveal the polarimetric signal with sufficient fidelity (Mierla et al., 2010; DeForest, de Koning, and Elliott, 2017). Many authors have more recently developed the theory and demonstrated the practice, as reviewed by Gibson et al. (2025). PUNCH is specifically designed to measure the 3D trajectory and structure of small ejecta, CMEs, and large solar-wind features such as stream interaction regions (SIRs).
@@ 2 PUNCH Science
Despite decades of study (e.g., Neugebauer and Snyder, 1966; Hundhausen, 1972; Lopez, 1987; Marsch and Tu, 1993; Sheeley et al., 1997a; Zurbuchen et al., 1998; Bruno et al., 2001; Borovsky, 2008; Suess et al., 2009; Cranmer, Gibson, and Riley, 2017; Viall and Borovsky, 2020; Viall, DeForest, and Kepko, 2021), quantitative insight into the origins of solar-wind structure and variability has been limited by a lack of direct observations. As a result, scientific studies have remained heavily reliant on modeling to connect the remotely observed corona with heliospheric in-situ data (Odstrcil and Pizzo, 2009; Provornikova et al., 2024). PUNCH is specifically designed to fill in this gap by imaging the evolving plasma with sufficient resolution and sensitivity to address the PUNCH science objectives, described in the subsections below.
@@ 2.1.1 Science Topic 1A: Solar Wind Flow
The Ulysses mission (McComas, Gosling, and Skoug, 2000; von Steiger et al., 2000; McComas et al., 2002) revolutionized understanding of the heliosphere when its measurements revealed the solar wind’s latitudinal structure. Being an in-situ mission, Ulysses sampled only a single narrow locus along each orbit, and required years to do so. In contrast, imagers reveal instantaneous global structure. Many authors have analyzed solar wind scceleration in the low corona (e.g., Sheeley et al., 1997a; Wang et al., 2000). Deep-field imaging of the mid to outer corona (DeForest et al., 2016, 2018) reveals faint moving structures and inhomogeneities in the outer corona at all latitudes and solar wind speeds (Figure 3), demonstrating the feasibility of tracing solar wind flow and its variability throughout the transition from coronal flow into the separated regime of the young solar wind. These measurements are necessary to constrain current incomplete understanding of solar wind acceleration flow, (e.g., Cranmer and Winebarger, 2019).
@@ 2.1.2 Science Topic 1B: Mesostructures and Turbulence
The solar wind near Earth is dominated by fluctuations of unknown origin. Mesoscale structures (those between the small-scale kinetic regime and large structures like coronal mass ejections or stream interaction regions) may form mainly from solar wind turbulence (Elliott et al., 2010; Matthaeus and Velli, 2011; DeForest et al., 2015), from solar-wind release (Sheeley et al., 1997a, 2009; Borovsky, 2008; Rouillard et al., 2010; Viall et al., 2010; Neugebauer, 2012; Hardwick et al., 2013; Kepko et al., 2016), or from a combination of both processes. The dichotomy of mechanisms, and open questions surrounding these mesostructures, [^4] The term “microstructure” has been used since at least the 1990s by the solar imaging community (Sheeley et al., 1997b). We use “mesostructures” instead, because the “micro-” prefix, used in the context of solar wind science, implies plasma kinetic scales – which are much smaller than the resolution of current imagers including PUNCH. were recently reviewed by Viall, DeForest, and Kepko (2021). Mesostructures are small in heliospheric imagers, but large compared to Earth’s magnetosphere; they are therefore important daily drivers of geospace dynamics and space weather (e.g., Viall, Kepko, and Spence, 2009; Di Matteo et al., 2022; Kurien et al., 2024).
@@
The characteristics and formation locations of mesostructures are important constraints on models of solar-wind origin and acceleration. In particular, they distinguish models that invoke steady flow along the edges of coronal holes (Suess et al., 1998; Vásquez, van Ballegooijen, and Raymond, 2003; Wang et al., 2007) from those that are intrinsically intermittent through interchange reconnection or multiple ejections of small plasmoids (Suess, Wang, and Wu, 1996; Fisk et al., 2003; Endeve, Holzer, and Leer, 2004; Rappazzo et al., 2005; Chen et al., 2009; Antiochos et al., 2011; Uritsky et al., 2021; Raouafi et al., 2023).
@@
Some mesostructures are present in the outer corona and visible in Thomson-scattered light. Two classes of mesoscale density features in the solar wind have been identified out to 0.3 AU; “puffs” (including quasiperiodic density features) that originate in the corona (Viall et al., 2010; Viall and Vourlidas, 2015), and “flocculae,” that become apparent at distances between 0.1 and 0.3 AU (DeForest et al., 2016). Both kinds of variability are faintly discernible with STEREO/SECCHI, but in-situ observations (Viall, Kepko, and Spence, 2008; Kepko, Viall, and DiMatteo, 2024) suggest a zoo of similar mesostructure at scales just below the sensitivity limit of that suite. The location and distribution of such features probes the $\beta=1$ surface, where the Sun’s magnetic field becomes too weak to stabilize flows against hydrodynamic instabilities in the solar wind.
@@
Evolution of solar wind mesostructures is also a strong probe into the development of large-scale turbulence in the young solar wind. Turbulence may be the key to solving the “solar-wind heating problem,” that the solar wind does not cool as expected from adiabatic expansion. For example, at 1 AU the temperatures are 100 times higher than expected from adiabatic propagation alone. It has been proposed that kinetic energy, entrained in large-scale turbulence, cascades downward to kinetic scales in the solar wind, heating it via plasma dissipative effects (Coleman, 1968; Leamon et al., 1998). In-situ measurements have revealed a turbulent cascade with a a well-established inertial range at 1 AU (Marsch and Tu, 1993; Goldstein and Roberts, 1999). Coronal observations suggest the existence of turbulence there as well (Coles and Harmon, 1989; Chae, Schühle, and Lemaire, 1998; Kohl et al., 2006; Tomczyk and McIntosh, 2009; Liu et al., 2014; Cranmer, 2020), but how and where a turbulent cascade develops remains unclear (Zank and Matthaeus, 1992; Chhiber et al., 2018). Analysis of images from STEREO/HI-1 (DeForest et al., 2016) has shown that the visible radial structures (“striae”) of the solar corona break up via currently unresolved processes at apparent distances between 10 and 20 (40 R to 80 R) from the Sun. The striae exhibit small lateral motions and spectral evolution, seen close to the HI-1 noise floor, that may indicate the development of a turbulent cascade (Figure 5).
@@ 2.1.3 Science Topic 1C: The Alfvén Surface
The coronal plasma is in a state of constant expansion into interplanetary space, where it forms the solar wind. The boundary where the radial flow speed exceeds the fast-mode magnetosonic wave speed is a critical surface beyond which information cannot propagate back down to the Sun, at least in the form of MHD waves (Weber and Davis, 1967; Hundhausen, 1972). This “Alfvén surface”, so called because the relevant fast-mode speed is the Alfvén speed, is a natural outer limit of the corona itself, marking a natural border between the extended solar corona and the disconnected solar wind. It is the true source surface of the interplanetary magnetic field, and its shape and location are strong discriminators among solar wind acceleration models. It also plays a role in stellar rotational spin-down (Garraffo, Drake, and Cohen, 2015), which is relevant to the Sun’s past and future evolution (Metcalfe, Egeland, and van Saders, 2016).
Models (e.g., Cranmer, van Ballegooijen, and Edgar, 2007; Verdini, Velli, and Buchlin, 2009; Zhao and Hoeksema, 2010; Goelzer, Schwadron, and Smith, 2014; Tasnim and Cairns, 2016) place the location of the Alfvén surface between about 10 and 30 R, with a complexity that evolves and changes with the solar magnetic cycle. Fourier filtering of images has revealed the existence of inbound features out to about 15 R (DeForest, Howard, and McComas, 2014), but there remain difficulties matching the observed speeds and accelerations with solar wind models (Verdini, Velli, and Buchlin, 2009; Tenerani, Velli, and DeForest, 2016). It may be the case that these inflowing parcels are associated with nonlinear MHD structures such as shocks, jets, or shear instabilities (Cranmer, DeForest, and Gibson, 2021), complementing expected outflowing parcels from reconnection in the outer corona and heliosphere (Gosling et al., 2005; Khabarova et al., 2016).
@@
PUNCH studies the location of the Alfvén zone through analysis to separate inbound and outbound features in the solar wind, and determine the region above which inbound motions cease. Across time, the shape, complexity, and evolution of this zone, and the Alfvén speed itself, are determined through direct measurement of feature motions. Although particular moving features may be due to bulk motion of mesostructures or to waves (Tenerani, Velli, and DeForest, 2016), in either case the Alfvén speed of the surrounding medium limits the relative speed of the bulk flow and the visible feature (Cranmer, DeForest, and Gibson, 2021), enabling PUNCH’s visual-tracking approach to identifying the Alfvén surface (or zone). PUNCH science on this topic is described in more detail by Cranmer et al. (2023).
@@ 2.2 Science Objective 2: The Dynamic Solar Wind
The impact of large, dynamic structures such as coronal mass ejections (CMEs), stream interaction regions (SIRs), and shocks on the solar wind and geospace is well documented (e.g., Schwenn, 2006; Pulkkinen, 2007). A deeper understanding of the physical processes controlling these space-weather drivers holds the key to improvements in predictive capability. However, tracking, monitoring, and diagnosing even the most basic physical properties of transients (speed, density, morphology) has remained limited and challenging. This motivates PUNCH’s improvements in resolution, FOV, and photometry, and also its polarimetric capabilities, which can improve predictions of CME and SIR location and arrival time (Howard et al., 2013). 3-D localization using polarimetry improves on unpolarized stereoscopy alone, by eliminating multiple-viewing-angle confusion effects. Large-scale structures may be located in several ways: constraint of kinematic envelope models (Howard et al., 2013), direct analytic location of substructures (DeForest, de Koning, and Elliott, 2017) and kinematically constrained tomographic inversion (Jackson et al., 2006; Bisi et al., 2008; Jackson et al., 2011; Morgan, 2015), as well as forward modeling (e.g., Talpeanu et al., 2020).
@@ 2.2.1 Science Topic 2A: Coronal Mass Ejections
Coronal mass ejections (CMEs) are the largest perturbations of the solar wind, and the source of the strongest space weather events at Earth. They have been objects of study since the 1970s. However, many aspects of CME physics, particularly how they propagate and evolve through the solar wind medium, have not been well constrained by measurement. Earlier imagers could track only the largest or brightest components of CME structure beyond 20 R, yielding limited information about how CME sub-structures interact with each other or with the solar wind, despite a handful of joint tracking/in-situ studies of particular events (e.g., Howard and DeForest, 2012a; DeForest, Howard, and McComas, 2013; Palmerio et al., 2024). PUNCH is designed to enable analysis of not only bulk CME bodies, including the front trajectory and possible bulk deflection, but also the interior features that may be remnants of the source structure in the lower corona.
The structure and orientation of CME magnetic fields at the Earth largely determine space-weather impact, representing a central challenge for heliophysics. Current understanding of CME magnetic structure is based on in-situ sampling (Gosling et al., 1987; Lepping, Jones, and Burlaga, 1990), in combination with near-Sun measurements of CMEs (Dere et al., 1999) and their precursors (Gibson, 2015). These observations generally support a model of CMEs as magnetic flux ropes (Vourlidas et al., 2013). However, CME magnetic structures are known to change in the solar wind due to distortion (Howard and DeForest, 2012b), deflection and rotation (Isavnin, Vourlidas, and Kilpua, 2014), CME-CME interactions (Gopalswamy et al., 2001), and magnetic reconnection/erosion (Gosling et al., 2005; Ruffenach et al., 2015). Assuming the axis of the CME flux rope can be obtained through model or bulk 3D measurement, its chirality follows from the direction of circulation about the axis (DeForest, de Koning, and Elliott, 2017).

### v2

@@ 1 Introduction
Joint studies of the solar corona and solar wind have generally required spanning the two radically different experimental disciplines, but have borne fruit in connecting the Sun and near-Earth environments (Zirker, 1977; Sheeley et al., 1985; Domingo et al., 1995). ESA’s Solar Orbiter mission (Müller et al., 2020) carries a hybrid instrument suite specifically designed to enable joint remote/in-situ studies of solar wind features and streams, by associating particular remotely sensed coronal features with the plasma detected in-situ by the spacecraft. Recently, the Parker Solar Probe (Fox et al., 2016) has carried in-situ sampling devices into the outer reaches of the solar corona itself (Kasper et al., 2021), allowing direct measurement of the microphysics in the outer corona with simultaneous remote sensing of the relevant large-scale structures surrounding the system under study.
@@
Multiple instruments, including the Coriolis Solar Mass Ejection Imager (Coriolis/SMEI; Eyles et al., 2003), the Solar-Terrestrial Relations Observatory’s Sun-Earth Connection Coronal and Heliospheric Investigation (STEREO/SECCHI; Howard et al., 2008), the Solar Orbiter Heliospheric Imager (SOLO/SOLOHI; Howard et al., 2020), and the Parker Solar Probe Wide-field Imager for Solar Probe (PSP/WISPR; Vourlidas et al., 2016b), have been built to image large and bright features in the solar wind. With the advent of imaging from off the Sun-Earth line, via STEREO/SECCHI, direct, quantitative, simultaneous imaging and in-situ studies of Earth-bound events became possible (e.g., Davies et al., 2009; DeForest et al., 2013a), and the presence of multiple deep-space missions with intersecting fields of view and on-board in-situ instruments has greatly increased the number of such studies. PUNCH brings routine global imaging of the solar wind and its structures into the mix of available observations, explicitly enabling a unified treatment of the corona and solar wind as a single system.
@@
PUNCH polarimetry is designed to allow location of features in three dimensions (3D) from a single vantage point: both mesoscale solar wind structure (e.g., Viall et al., 2021) and coronal mass ejections (de Koning et al., 2009; Mierla et al., 2009; DeForest et al., 2017). This is possible because the K corona consists of sunlight that is Thomson scattered by free electrons. Thomson scattering polarizes the outbound beam according to the scattering angle (Billings, 1966; Tappin and Howard, 2009; DeForest et al., 2013b). Each coronal feature is therefore polarized to a degree that depends on its three-dimensional location. The basic theory has been known for decades, but was only reduced to practice once image quality and analysis methods improved enough to reveal the polarimetric signal with sufficient fidelity (Mierla et al., 2010; DeForest et al., 2017). Many authors have more recently developed the theory and demonstrated the practice, as reviewed by Gibson et al. (2025). PUNCH is specifically designed to measure the 3D trajectory and structure of small ejecta, CMEs, and large solar-wind features such as stream interaction regions (SIRs).
@@ 2 PUNCH Science
Despite decades of study (e.g., Neugebauer and Snyder, 1966; Hundhausen, 1972; Lopez, 1987; Marsch and Tu, 1993; Sheeley et al., 1997a; Zurbuchen et al., 1998; Bruno et al., 2001; Borovsky, 2008; Suess et al., 2009; Cranmer et al., 2017; Viall and Borovsky, 2020; Viall et al., 2021), quantitative insight into the origins of solar-wind structure and variability has been limited by a lack of direct observations. As a result, scientific studies have remained heavily reliant on modeling to connect the remotely observed corona with heliospheric in-situ data (Odstrcil and Pizzo, 2009; Provornikova et al., 2024). PUNCH is specifically designed to fill in this gap by imaging the evolving plasma with sufficient resolution and sensitivity to address the PUNCH science objectives, described in the subsections below.
@@ 2.1.1 Science Topic 1A: Solar Wind Flow
The Ulysses mission (McComas et al., 2000; von Steiger et al., 2000; McComas et al., 2002) revolutionized understanding of the heliosphere when its measurements revealed the solar wind’s latitudinal structure. Being an in-situ mission, Ulysses sampled only a single narrow locus along each orbit, and required years to do so. In contrast, imagers reveal instantaneous global structure. Many authors have analyzed solar wind scceleration in the low corona (e.g., Sheeley et al., 1997a; Wang et al., 2000). Deep-field imaging of the mid to outer corona (DeForest et al., 2016, 2018) reveals faint moving structures and inhomogeneities in the outer corona at all latitudes and solar wind speeds (Figure 3), demonstrating the feasibility of tracing solar wind flow and its variability throughout the transition from coronal flow into the separated regime of the young solar wind. These measurements are necessary to constrain current incomplete understanding of solar wind acceleration flow, (e.g., Cranmer and Winebarger, 2019).
@@ 2.1.2 Science Topic 1B: Mesostructures and Turbulence
The solar wind near Earth is dominated by fluctuations of unknown origin. Mesoscale structures (those between the small-scale kinetic regime and large structures like coronal mass ejections or stream interaction regions) may form mainly from solar wind turbulence (Elliott et al., 2010; Matthaeus and Velli, 2011; DeForest et al., 2015), from solar-wind release (Sheeley et al., 1997a, 2009; Borovsky, 2008; Rouillard et al., 2010; Viall et al., 2010; Neugebauer, 2012; Hardwick et al., 2013; Kepko et al., 2016), or from a combination of both processes. The dichotomy of mechanisms, and open questions surrounding these mesostructures, [^4] The term “microstructure” has been used since at least the 1990s by the solar imaging community (Sheeley et al., 1997b). We use “mesostructures” instead, because the “micro-” prefix, used in the context of solar wind science, implies plasma kinetic scales – which are much smaller than the resolution of current imagers including PUNCH. were recently reviewed by Viall et al. (2021). Mesostructures are small in heliospheric imagers, but large compared to Earth’s magnetosphere; they are therefore important daily drivers of geospace dynamics and space weather (e.g., Viall et al., 2009; Di Matteo et al., 2022; Kurien et al., 2024).
@@
The characteristics and formation locations of mesostructures are important constraints on models of solar-wind origin and acceleration. In particular, they distinguish models that invoke steady flow along the edges of coronal holes (Suess et al., 1998; Vásquez et al., 2003; Wang et al., 2007) from those that are intrinsically intermittent through interchange reconnection or multiple ejections of small plasmoids (Suess et al., 1996; Fisk et al., 2003; Endeve et al., 2004; Rappazzo et al., 2005; Chen et al., 2009; Antiochos et al., 2011; Uritsky et al., 2021; Raouafi et al., 2023).
@@
Some mesostructures are present in the outer corona and visible in Thomson-scattered light. Two classes of mesoscale density features in the solar wind have been identified out to 0.3 AU; “puffs” (including quasiperiodic density features) that originate in the corona (Viall et al., 2010; Viall and Vourlidas, 2015), and “flocculae,” that become apparent at distances between 0.1 and 0.3 AU (DeForest et al., 2016). Both kinds of variability are faintly discernible with STEREO/SECCHI, but in-situ observations (Viall et al., 2008; Kepko et al., 2024) suggest a zoo of similar mesostructure at scales just below the sensitivity limit of that suite. The location and distribution of such features probes the $\beta=1$ surface, where the Sun’s magnetic field becomes too weak to stabilize flows against hydrodynamic instabilities in the solar wind.
@@
Evolution of solar wind mesostructures is also a strong probe into the development of large-scale turbulence in the young solar wind. Turbulence may be the key to solving the “solar-wind heating problem,” that the solar wind does not cool as expected from adiabatic expansion. For example, at 1 AU the temperatures are 100 times higher than expected from adiabatic propagation alone. It has been proposed that kinetic energy, entrained in large-scale turbulence, cascades downward to kinetic scales in the solar wind, heating it via plasma dissipative effects (Coleman, 1968; Leamon et al., 1998). In-situ measurements have revealed a turbulent cascade with a a well-established inertial range at 1 AU (Marsch and Tu, 1993; Goldstein and Roberts, 1999). Coronal observations suggest the existence of turbulence there as well (Coles and Harmon, 1989; Chae et al., 1998; Kohl et al., 2006; Tomczyk and McIntosh, 2009; Liu et al., 2014; Cranmer, 2020), but how and where a turbulent cascade develops remains unclear (Zank and Matthaeus, 1992; Chhiber et al., 2018). Analysis of images from STEREO/HI-1 (DeForest et al., 2016) has shown that the visible radial structures (“striae”) of the solar corona break up via currently unresolved processes at apparent distances between 10 and 20 (40 R to 80 R) from the Sun. The striae exhibit small lateral motions and spectral evolution, seen close to the HI-1 noise floor, that may indicate the development of a turbulent cascade (Figure 5).
@@ 2.1.3 Science Topic 1C: The Alfvén Surface
The coronal plasma is in a state of constant expansion into interplanetary space, where it forms the solar wind. The boundary where the radial flow speed exceeds the fast-mode magnetosonic wave speed is a critical surface beyond which information cannot propagate back down to the Sun, at least in the form of MHD waves (Weber and Davis, 1967; Hundhausen, 1972). This “Alfvén surface”, so called because the relevant fast-mode speed is the Alfvén speed, is a natural outer limit of the corona itself, marking a natural border between the extended solar corona and the disconnected solar wind. It is the true source surface of the interplanetary magnetic field, and its shape and location are strong discriminators among solar wind acceleration models. It also plays a role in stellar rotational spin-down (Garraffo et al., 2015), which is relevant to the Sun’s past and future evolution (Metcalfe et al., 2016).
Models (e.g., Cranmer et al., 2007; Verdini et al., 2009; Zhao and Hoeksema, 2010; Goelzer et al., 2014; Tasnim and Cairns, 2016) place the location of the Alfvén surface between about 10 and 30 R, with a complexity that evolves and changes with the solar magnetic cycle. Fourier filtering of images has revealed the existence of inbound features out to about 15 R (DeForest et al., 2014), but there remain difficulties matching the observed speeds and accelerations with solar wind models (Verdini et al., 2009; Tenerani et al., 2016). It may be the case that these inflowing parcels are associated with nonlinear MHD structures such as shocks, jets, or shear instabilities (Cranmer et al., 2021), complementing expected outflowing parcels from reconnection in the outer corona and heliosphere (Gosling et al., 2005; Khabarova et al., 2016).
@@
PUNCH studies the location of the Alfvén zone through analysis to separate inbound and outbound features in the solar wind, and determine the region above which inbound motions cease. Across time, the shape, complexity, and evolution of this zone, and the Alfvén speed itself, are determined through direct measurement of feature motions. Although particular moving features may be due to bulk motion of mesostructures or to waves (Tenerani et al., 2016), in either case the Alfvén speed of the surrounding medium limits the relative speed of the bulk flow and the visible feature (Cranmer et al., 2021), enabling PUNCH’s visual-tracking approach to identifying the Alfvén surface (or zone). PUNCH science on this topic is described in more detail by Cranmer et al. (2023).
@@ 2.2 Science Objective 2: The Dynamic Solar Wind
The impact of large, dynamic structures such as coronal mass ejections (CMEs), stream interaction regions (SIRs), and shocks on the solar wind and geospace is well documented (e.g., Schwenn, 2006; Pulkkinen, 2007). A deeper understanding of the physical processes controlling these space-weather drivers holds the key to improvements in predictive capability. However, tracking, monitoring, and diagnosing even the most basic physical properties of transients (speed, density, morphology) has remained limited and challenging. This motivates PUNCH’s improvements in resolution, FOV, and photometry, and also its polarimetric capabilities, which can improve predictions of CME and SIR location and arrival time (Howard et al., 2013). 3-D localization using polarimetry improves on unpolarized stereoscopy alone, by eliminating multiple-viewing-angle confusion effects. Large-scale structures may be located in several ways: constraint of kinematic envelope models (Howard et al., 2013), direct analytic location of substructures (DeForest et al., 2017) and kinematically constrained tomographic inversion (Jackson et al., 2006; Bisi et al., 2008; Jackson et al., 2011; Morgan, 2015), as well as forward modeling (e.g., Talpeanu et al., 2020).
@@ 2.2.1 Science Topic 2A: Coronal Mass Ejections
Coronal mass ejections (CMEs) are the largest perturbations of the solar wind, and the source of the strongest space weather events at Earth. They have been objects of study since the 1970s. However, many aspects of CME physics, particularly how they propagate and evolve through the solar wind medium, have not been well constrained by measurement. Earlier imagers could track only the largest or brightest components of CME structure beyond 20 R, yielding limited information about how CME sub-structures interact with each other or with the solar wind, despite a handful of joint tracking/in-situ studies of particular events (e.g., Howard and DeForest, 2012a; DeForest et al., 2013a; Palmerio et al., 2024). PUNCH is designed to enable analysis of not only bulk CME bodies, including the front trajectory and possible bulk deflection, but also the interior features that may be remnants of the source structure in the lower corona.
The structure and orientation of CME magnetic fields at the Earth largely determine space-weather impact, representing a central challenge for heliophysics. Current understanding of CME magnetic structure is based on in-situ sampling (Gosling et al., 1987; Lepping et al., 1990), in combination with near-Sun measurements of CMEs (Dere et al., 1999) and their precursors (Gibson, 2015). These observations generally support a model of CMEs as magnetic flux ropes (Vourlidas et al., 2013). However, CME magnetic structures are known to change in the solar wind due to distortion (Howard and DeForest, 2012b), deflection and rotation (Isavnin et al., 2014), CME-CME interactions (Gopalswamy et al., 2001), and magnetic reconnection/erosion (Gosling et al., 2005; Ruffenach et al., 2015). Assuming the axis of the CME flux rope can be obtained through model or bulk 3D measurement, its chirality follows from the direction of circulation about the axis (DeForest et al., 2017).